OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -6, an = 4 • an-1 0, 4, -24, -20, -16, -12 -24, -96, -384, -1536, -6144, -24,576 -6, -24, -20, -16, -12, -8 -6, -24, -96, -384, -1536, -6144

OpenStudy (anonymous):

i know its either c or d ..... def. not a or b because it clearly tells you a1 = -6

OpenStudy (anonymous):

@akonkel @idku

OpenStudy (tylerd):

d

OpenStudy (tylerd):

an=4*an-1 right? that means -24=-6*4 -96=-24*4 etc etc

OpenStudy (anonymous):

yeah okay i see that now.. thank you (: wanna help with another?

OpenStudy (tylerd):

sure

OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n3 + 2n is divisible by 3 for every positive integer n.

OpenStudy (tylerd):

is that n^3 or 3n?

OpenStudy (anonymous):

n^3

OpenStudy (anonymous):

( n + 1 )^3 + 2( n + 1 ) = ( n^3 + 3n^2 + 3n + 1 ) + ( 2n + 2 ) = ( n^3 + 2n ) + ( 3n^2 + 3n + 3 ) = ( n^3 + 2n ) + 3( n^2 + n + 1 ) Therefore ( n^3 + 2n ) is divisible by 3 right?

OpenStudy (tylerd):

im actually not familiar with this sorry.

OpenStudy (anonymous):

what about this one ? Write the sum using summation notation, assuming the suggested pattern continues. 64 + 81 + 100 + 121 + ... + n2 + ... summation of n squared from n equals eight to infinity summation of n minus one squared from n equals eight to infinity summation of n squared from n equals nine to infinity summation of n plus one squared from n equals eight to infinity

OpenStudy (anonymous):

@TylerD

OpenStudy (tylerd):

looks like the first one

OpenStudy (tylerd):

n^2 starting at n=8 so it starts at 8^2 which = 64 then the next term would be 9^2 = 81 next term 10^2 = 100 etc etc etc